# Generating Fine-Grained Human Motions Using ChatGPT-Refined Descriptions

### v1

@@ Abstract
Recently, significant progress has been made in text-based motion generation, enabling the generation of diverse and high-quality human motions that conform to textual descriptions. However, it remains challenging to generate fine-grained or stylized motions due to the lack of datasets annotated with detailed textual descriptions. By adopting a divide-and-conquer strategy, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for human motion generation. Specifically, we first parse previous vague textual annotation into fine-grained description of different body parts by leveraging a large language model (GPT-3.5). We then use these fine-grained descriptions to guide a transformer-based diffusion model. FG-MDM can generate fine-grained and stylized motions even outside of the distribution of the training data. Our experimental results demonstrate the superiority of FG-MDM over previous methods, especially the strong generalization capability. We will release our fine-grained textual annotations for HumanML3D and KIT.
Figure 1: FG-MDM can generate high-quality fine-grained and stylized motions by using fine-grained descriptions of different body parts.
@@ 1 Introduction
Human motion generation is an important research topic in communities of both computer vision and computer graphics. It aims to simulate and generate realistic human movements using computers. With the advancement of technologies such as virtual reality, augmented reality, and movie special effects, there is a growing demand for high-quality human motion generation. In recent years, quite a few innovative methods and techniques have emerged to tackle this challenging task [45]. Deep generative models, including GANs [20, 1], VAEs [9, 25, 24], and Diffusion models [35, 42, 16, 5], have been widely applied to human motion generation.
However, there is relatively less research on generating fine-grained motions, which usually mobilize the entire body and involve multiple actions, based on textual descriptions. In [3], Athanasiou et al. proposed a spatial composition framework named SINC for simultaneous action generation. It should be noted that SINC was designed for motion generation based on action labels, not using natural language. MotionDiffuse can achieve fine-grained controlling when generating motions using textual descriptions [42]. However, the controlling ability is largely limited by the scarcity of the training data. Likewise, there are very few research efforts devoted to stylized motion generation with natural language, which generates motions accompanied with rich styles and emotions. In [34], Tevet et al. proposed MotionCLIP to study stylized motion generation. However, only simple action classes with several style labels were considered. Stylized motions defined by textual descriptions were not investigated.
@@
A challenging issue of current human motion generation is the scarcity of datasets. This is especially true for datasets that include fine-grained descriptions or stylized text annotations. Datasets, such as KIT [26], BABEL [27], and HumanML3D [9], provide rich motion information. Their textual descriptions instead are often simple and direct, lacking stylized and fine-grained descriptions. Therefore, compared to traditional human motion generation, generating fine-grained and stylized motions is more challenging due to the limited scale and diversity of existing motion capture datasets. In Figure 3 and Figure 4, we show some fine-grained and stylized motions generated by MDM [35] and MLD [5]. Although both approaches achieve impressive performance in human motion generation, they show poor generalization ability since these fine-grained and stylized motions are outside of the distribution of the training data. Recently, there have been attempts to leverage human mesh recovery approaches [36, 40, 39] to collect large-scale text pseudo-pose datasets [4]. As shown in Azadi et al. [4], the pre-training on such datasets can significantly improve the generalization to in-the-wild descriptions, but generating fine-grained and stylized motions remains challenging due to the lack of detailed descriptions.
@@
Then, with only limited motion capture datasets available, can we still achieve fine-grained and stylized motion generation? For a textual description defining a fine-grained or stylized motion, is there any way to associate it with motions within the dataset? Note that the fine-grained and stylized motions might be outside of the distribution of the dataset. For a never-before-seen motion, the motion of the entire body is indeed unseen, but motions of specific body parts may be inside of the distribution of the dataset. Therefore, we can adopt a divide-and-conquer strategy. By re-annotating the motion for different body parts with fine-grained descriptions, we can associate these body parts with specific body parts within the dataset. For example, the stylized description “A person walks depressingly.” can be reformulated as “His arms hang heavily by his sides. His legs move slowly, taking short steps with little energy…”. Leg movement in this stylized motion may appear in “A person walks aimlessly and slowly.”. The arm movement may appear in “A person stands in a relaxed manner.”. In Figure LABEL:figshow, we give two examples. On one hand, adopting fine-grained descriptions allows the model to get how the fine-grained or stylized motion is performed in detail. On the other hand, re-annotating the motion for different body parts enables the model to handle motions beyond the scope of the dataset.
@@
Although manually annotating fine-grained textual descriptions for body parts provides more accurate data, it requires a significant amount of manual work, resulting in huge cost. Fortunately, with the rapid development of large language models, OpenAI’s GPT series models [23], known for their outstanding natural language processing capabilities, have gained widespread attention worldwide. In [7], Gilardi et al. demonstrated that ChatGPT performs as well as human annotators in some tasks. In [15], Action-GPT explores the excellent capability of ChatGPT in expanding human action descriptions. However, the generated content tends to be excessively redundant. For our task, we carefully design a prompt that allows ChatGPT-3.5 to provide detailed but non-redundant transcriptions of text descriptions about the human motion. We then use this prompt and ChatGPT-3.5 to transcribe 44,970 short text descriptions from HumanML3D [9] and 6,353 text descriptions from KIT [26] for fine-grained and stylized motion generation. We will make these fine-grained transcriptions publicly available.
@@
With these fine-grained descriptions, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for human motion generation. Specifically, we replace the original simple and direct text with ChatGPT-generated fine-grained descriptions of individual body parts to guide a transformer-based diffusion model. Following MDM [35], we encode the entire fine-grained description with CLIP [29] as a global token of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens. By adopting these two tokens, the model can pay attention to both the global and detailed information of human motions, thereby improving the accuracy and completeness of the denoising results.
@@
  • We present a novel framework that utilizes fine-grained descriptions of different body parts to guide the training of the human motion diffusion model. This framework enables the generation of fine-grained and stylized motions even in the absence of datasets annotated with fine-grained and stylized descriptions.
  • We carefully design a prompt that enables ChatGPT to convert short texts into detailed but non-redundant descriptions of different body parts. We then use this prompt to transcribe 44,970 texts from HumanML3D and 6,353 texts from KIT into fine-grained descriptions. We will make these fine-grained transcriptions publicly available.
  • We conduct a series of experiments to evaluate our model’s ability in not only fitting the training data but also generating fine-grained and stylized motions, i.e., the generalization capability.
@@ 2.1 Human Motion Generation
Recent years have seen a great interest in human motion generation. Previous work has explored unconditional generative models [44, 28] as well as generative models using various input conditions, such as text [9, 35, 5], prior motion [22, 10], action class [8, 24], and music [18, 37, 17]. In this paper, we focus on text-to-motion generation. Early work usually addressed the text-to-motion task with a sequence-to-sequence model [19]. Later on, the focus shifted beyond simple action labels. For example, Guo et al. utilized variational autoencoders to generate motions from text [9] , significantly enhancing the quality and diversity of generated motions. With the success of diffusion models in AIGC, MDM [35] and other related work [42, 16, 5] have introduced diffusion models into the text-to-motion domain, resulting in impressive achievements.
Recently, there has been some work attempting to generate fine-grained or stylized motions. For example, MotionCLIP [34] trains an autoencoder to align motion with the text and images of CLIP in the latent feature space, exploring the generation of stylized actions. MotionDiffuse [42] instead proposes precise control of motions and stylization of body parts using noise interpolation based on the properties of diffusion models. GestureDiffuCLIP [2] incorporates style guidance into the diffusion model through AdaIN [13], enabling the generation of realistic and stylized gestures. For music-to-motion, there has been work [18, 37] trying to generate dance movements of different styles by using different styles of music.
The work most closely related to ours is Action-GPT [15], which introduced, for the first time, Large Language Models (LLMs) into the field of text-conditioned motion generation. Action-GPT can be integrated into any text-to-motion model. However, it enriches only the description of action classes, without providing detailed descriptions of different body parts and further guiding the model training accordingly. For simultaneous action generation, SINC [3] incorporates ChatGPT to identify the body parts involved in the textual description. By generating multiple motions and concatenating them using different body parts, it achieves impressive results. Specifically, SINC divides the human body into [’left arm’, ’right arm’, ’left leg’, ’right leg’, ’torso’, ’neck’, ’buttocks’, ’waist’], which we borrow from in our work. By fine-tuning LLMs, MotionGPT [14, 43] designs a pre-trained motion language model which supports various motion-related tasks through prompts.
@@ 2.2 Diffusion Generative Models
The diffusion model is a type of neural generative model based on the stochastic diffusion process in thermodynamics [33, 12]. It starts with samples from the data distribution and gradually adds noise through a forward diffusion process. Then, a neural network learns the reverse process to progressively remove the noise and restore the samples to their original states. Diffusion generative models have achieved significant success in image generation field [30, 31]. For conditional generation,  [6] introduced classifier-guided diffusion, while  [11] proposed a classifier-free method. Given their excellent generation quality,  [35, 42, 16] incorporated diffusion models into motion generation domain, leading to impressive results.
@@ 3 Method
Given a textual description, our goal is to generate a human motion $x^{1:n}=\{x^{i}\}_{i=1}^{n}$ that matches the given description. The motion consists of $n$ frames of human poses. For each pose $x^{i}\in\mathbb{R}^{J\times D}$, we represent it by joint rotations or positions, where $J$ represents the number of joints and $D$ represents the dimensionality of the joint representation. In Figure 2, we give an overview of our fine-grained human motion diffusion model. First, we adopt ChatGPT to perform fine-grained paraphrasing of the given vague textual description. This expands concise textual descriptions into descriptions of different body parts. FG-MDM then uses these fine-grained descriptions to guide a diffusion model for human motion generation.
@@ 3.1 Prompt Strategy
We first introduce the prompt strategy adopted for generating fine-grained descriptions. For a given textual description of a motion, we utilize ChatGPT-3.5 to generate more fine-grained descriptions based on different body parts. ChatGPT is a conversational model based on a large language model that can engage in natural language conversations and generate corresponding responses. The answers from ChatGPT are often directly influenced by the information and expression provided in the prompt. If the prompt offers clear and detailed questions or instructions, ChatGPT can typically provide relevant and accurate answers. However, if the prompt is too simple, ambiguous, or unclear, ChatGPT may generate unexpected responses or express unclear content. For our task, we carefully design an effective prompt by using experimental verification.
Our designed strategy is: “Translate the motion described by the given sentences to the motion of each body part only using one paragraph. The available body parts include [‘arms’, ‘legs’, ‘torso’, ‘neck’, ‘buttocks’, ‘waist’]. Here are some examples:[Q…A…]. Question: [sentence]”.
@@ 3.2 Diffusion Model for Motion Generation
The diffusion model consists of two parts: the forward process and the reverse process, both of which follow the Markov chain. The forward process involves adding noise. The input is the original motion $x_{0}^{1:n}$ from the data distribution, and the output is the motion $x_{t}^{1:n}$ with $t$ times of Gaussian noise added. When enough noise is added, the motion $x_{T}^{1:n}$ can approach the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, aims to reduce the noise in the Gaussian noise $x_{T}^{1:n}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In the denoising process, at diffusion step $t$, a portion of the noise is eliminated, resulting in a less noisy motion $x_{t-1}^{1:n}$. This step is repeated iteratively until the noise is completely removed, generating the denoised motion $\hat{x}_{0}^{1:n}$.
@@ 3.2.1 Network
we adopt a simple transformer [38] encoder architecture to implement our network $G$. Unlike the conventional diffusion model mentioned above, we follow [30] and predict the clean motion $\hat{x}_{0}^{1:n}$ instead of predicting the noise added at each diffusion time-step. The input of $G$ is the noised motion $x_{t}^{1:n}$ obtained by adding noise $t$ times to the original motion $x_{0}^{1:n}$. The noised motion $x_{t}^{1:n}$, together with the text condition code $GL$, $PT_{1:k}$ and the noise time-step $t$, is inputted to the transformer encoder, resulting in the clean motion $\hat{x}_{0}^{1:n}$. One of the reasons for directly predicting the clean motion in each step of the diffusion model is to incorporate human geometric losses during the training of the network, making the generated human motions more natural. At each sampling step $t$, from $T$ to 1, our model predicts the clean motion $\hat{x}_{0}^{1:n}$, and then adds noise back to $x_{t-1}^{1:n}$. After $T$ iterations, the final clean motion $\hat{x}_{0}^{1:n}$ is obtained.

### v2

# FG-MDM: Towards Zero-Shot Human Motion Generation via Fine-Grained Descriptions

@@ Abstract
Recently, significant progress has been made in text-based motion generation, enabling the generation of diverse and high-quality human motions that conform to textual descriptions. However, generating motions beyond the distribution of original datasets remains challenging, i.e., zero-shot generation. By adopting a divide-and-conquer strategy, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for zero-shot human motion generation. Specifically, we first parse previous vague textual annotations into fine-grained descriptions of different body parts by leveraging a large language model. We then use these fine-grained descriptions to guide a transformer-based diffusion model, which further adopts a design of part tokens. FG-MDM can generate human motions beyond the scope of original datasets owing to descriptions that are closer to motion essence. Our experimental results demonstrate the superiority of FG-MDM over previous methods in zero-shot settings. We will release our fine-grained textual annotations for HumanML3D and KIT.
Figure 1: FG-MDM can generate high-quality human motions in zero-shot settings by using fine-grained descriptions of different body parts.
@@ 1 Introduction
Human motion generation is an important research topic in communities of both computer vision and computer graphics. It aims to simulate and generate realistic human movements using computers. With the advancement of technologies such as virtual reality, augmented reality, and movie special effects, there is a growing demand for high-quality human motion generation. In recent years, several innovative methods and techniques have emerged to tackle this challenging task [52]. Deep generative models, including GANs [24, 1], VAEs [9, 31, 29], and diffusion models [41, 48, 18, 6], have been widely applied to human motion generation.
However, there is relatively less research on generating motions in zero-shot settings. In [40], Tevet et al. proposed MotionCLIP to align human motions with the CLIP space, implicitly injecting the rich semantic knowledge from CLIP into the motion domain to enhance zero-shot generation capability. AvatarCLIP [14] also utilized CLIP to implement a zero-shot text-driven framework for 3D avatar generation and animation. Liang et al.[53] pre-trained a large-scale unconditional diffusion model to learn rich out-of-domain motion traits. A Motion ControlNet with a Mixture-of-Controllers block was proposed to align CLIP token embeddings with motion features during fine-tuning.
@@
In order to improve the generalization capability of motion generation models, there have been also attempts to leverage human mesh recovery approaches [42, 45, 46] to collect large-scale pseudo-text-pose datasets [4]. As shown in Azadi et al. [4] and Lin et al. [23], the pre-training on such text-pose datasets can improve the generalization to in-the-wild descriptions. However, the static nature of text-pose data makes it difficult to well represent dynamic motions. Compared to traditional human motion generation, generating motions beyond the distribution of the dataset is more challenging due to the limited scale and diversity of existing motion capture datasets. In Figure 3 and Figure 4, we show some motions generated by MDM [41] and MLD [6]. Although both approaches achieve impressive performance in human motion generation, they show poor generalization ability since these motions are outside the training data distribution.
@@
Then, with only limited motion capture datasets available, can we still generate motions beyond the distribution of the dataset? For a textual description defining a motion beyond the distribution of the dataset, is there any way to associate it with motions within the dataset? For a never-before-seen motion, the entire body’s motion is indeed unseen, but motions of specific body parts might be inside the dataset. Therefore, we can adopt a divide-and-conquer strategy. By re-annotating the motion for different body parts with fine-grained descriptions, we can associate these body parts with specific body parts within the dataset. For example, a vague description “A person walks depressingly.” can be reformulated as “His arms hang heavily by his sides. His legs move slowly, taking short steps with little energy…”. Leg movement in this vague unseen motion may appear in “A person walks aimlessly and slowly.”, of which the motion is included in the dataset. And the arm movement may appear in “His arms hang heavily by his sides.”, of which the motion is included in the dataset. We give two examples in Figure 1. On one hand, adopting fine-grained descriptions allows the model to understand how the unseen motions are performed in detail. On the other hand, re-annotating motions for different body parts enables the model to learn the essence of motions better. Using fine-grained textual descriptions, we aim to improve the model’s zero-shot understanding capability.
@@
Although annotating fine-grained textual descriptions manually for body parts provides more accurate data, it requires significant manual work, resulting in huge costs. Fortunately, with the rapid development of large language models, OpenAI’s GPT series models [28], known for their outstanding natural language processing capabilities, have gained widespread attention worldwide. In [8], Gilardi et al. demonstrated that ChatGPT performs as well as human annotators in some tasks. In [17], Action-GPT explores the excellent capability of ChatGPT in expanding human action descriptions. However, the generated content tends to be excessively redundant. For our task, we carefully design a prompt that allows ChatGPT-3.5 to provide detailed but non-redundant transcriptions of text descriptions about human motion. We then use this prompt and ChatGPT-3.5 to transcribe 44,970 short text descriptions from HumanML3D [9] and 6,353 text descriptions from KIT [33] for model training. We will make these fine-grained transcriptions available to the public.
@@
With these fine-grained descriptions, we propose a new framework named Fine-Grained Human Motion Diffusion Model (FG-MDM) for human motion generation. Specifically, we replace the original simple and vague text with ChatGPT-generated fine-grained descriptions of individual body parts to guide a transformer-based diffusion model. Following MDM [41], we encode the entire fine-grained description with CLIP [35] as a global token of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens. By adopting these tokens, the model can pay attention to both the global and detailed information of human motions, thereby improving the accuracy and completeness of the denoising results.
@@
  • We present a novel framework that utilizes fine-grained descriptions of different body parts to guide the denoising process of the transformer-based diffusion model. This framework is capable of generating a broader range of motions that extend beyond the distribution of training datasets.
  • We carefully design a prompt that enables ChatGPT to convert short and vague texts into detailed but non-redundant descriptions of different body parts. We then use this prompt to transcribe 44,970 texts from HumanML3D and 6,353 texts from KIT into fine-grained descriptions. We will make these fine-grained transcriptions publicly available.
  • We conduct a series of experiments to evaluate our model’s ability to not only fit the training data but also generate motions beyond the distribution of the dataset, i.e., the generalization capability.
@@ 2.1 Human Motion Generation
There has been a great interest in human motion generation in recent years. Previous work has explored unconditional generative models [51, 34] as well as generative models using various input conditions, such as text [9, 41, 6], prior motion [27, 11], action class [10, 29], and music [20, 43, 19]. In this paper, we focus on text-to-motion generation. Early work usually addressed the text-to-motion task with a sequence-to-sequence model [21]. Later on, the focus shifted beyond simple action labels. For example, Guo et al. utilized variational autoencoders to generate motions from text [9], significantly enhancing the quality and diversity of generated motions. With the success of diffusion models in AIGC, MDM [41] and other related work [48, 18, 6] have introduced diffusion models into the text-to-motion domain, resulting in impressive achievements.
There is a relative scarcity of work that directly focuses on the zero-shot capabilities of motion generation models. However, there has been some work attempting to generate fine-grained or stylized motions to expand the distribution of generated motions. For example, MotionCLIP [40] trained an autoencoder to align motion with the text and images of CLIP in the latent feature space, exploring the generation of unseen motions. MotionDiffuse [48] instead proposed precise control of motions and stylization of body parts using noise interpolation based on the properties of diffusion models. GestureDiffuCLIP [2] incorporated style guidance into the diffusion model through AdaIN [15], generating realistic and stylized gestures. FineMoGen [49] adopts a novel transformer architecture for generating and editing fine-grained motions. However, there is a significant difference between the methods above and ours. They overlooked that the essence of complex motions described by vague text is the combination of simple movements of body parts. This renders them ineffective when confronted with descriptions of motions they have not encountered.
The work most closely related to ours is Action-GPT [17], which introduced, for the first time, large language models into the field of text-conditioned motion generation. Action-GPT can be integrated into any text-to-motion model. However, it enriched only the description of action classes without providing detailed descriptions of different body parts and guiding the model training. For another action generation model, SINC [3] incorporated ChatGPT to identify the body parts involved in the textual description. It achieved impressive results by generating multiple motions and concatenating them using different body parts. Specifically, SINC divides the human body into [‘left arm’, ‘right arm’, ‘left leg’, ‘right leg’, ‘torso’, ‘neck’, ‘buttocks’, ‘waist’], which we borrow from in our work. There is also a costly method that utilizes LLMs. By fine-tuning LLMs, MotionGPT [16, 50] designed a pre-trained motion language model that supports various motion-related tasks through prompts. In contrast, our method is more efficient and can rapidly enhance the model’s zero-shot generation capabilities.
@@ 2.2 Diffusion Generative Models
The diffusion model is a neural generative model based on the stochastic diffusion process in thermodynamics [39, 12]. It starts with samples from the data distribution and gradually adds noise through a forward diffusion process. Then, a neural network learns the reverse process to progressively remove the noise and restore the samples to their original states. Diffusion generative models have achieved significant success in the image generation field [36, 37]. For conditional generation,  [7] introduced classifier-guided diffusion, while  [13] proposed a classifier-free method. Given their excellent generation quality, [41, 48, 18] incorporated diffusion models into the motion generation domain, leading to impressive results.
@@ 3 Method
Given a textual description, our goal is to generate a human motion $x^{1:n}=\{x^{i}\}_{i=1}^{n}$ that matches the given description. The motion consists of $n$ frames of human poses. For each pose $x^{i}\in\mathbb{R}^{J\times D}$, we represent it by joint rotations or positions, where $J$ represents the number of joints and $D$ represents the dimensionality of the joint representation. In Figure 2, we give an overview of our fine-grained human motion diffusion model. First, we adopt ChatGPT to perform fine-grained paraphrasing of the vague textual description. This expands concise textual descriptions into descriptions of different body parts. FG-MDM then uses these fine-grained descriptions to guide a diffusion model for human motion generation.
@@ 3.1 Prompt Strategy
We first introduce the prompt strategy adopted for generating fine-grained descriptions. We utilize ChatGPT-3.5 to create more fine-grained descriptions based on different body parts for a given textual description of a motion. ChatGPT is a conversational model based on a large language model that can engage in natural conversations and generate corresponding responses. The answers from ChatGPT are often directly influenced by the information and expression provided in the prompt. If the prompt offers clear and detailed questions or instructions, ChatGPT can typically provide relevant and accurate answers. However, if the prompt is too simple, ambiguous, or unclear, ChatGPT may generate unexpected responses or express unclear content. For our task, we carefully design an effective prompt by using experimental verification.
Our designed prompt is: “Translate the motion described by the given sentences to the motion of each body part only using one paragraph. The available body parts include [‘arms’, ‘legs’, ‘torso’, ‘neck’, ‘buttocks’, ‘waist’]. Here are some examples:[Q…A…]. Question: [sentence]”. [sentence] is the vague textual description that needs to be refined. [Q…A…] are four examples of Q&A pairs designed manually. Here is one of the examples,
Question: A man kicks something or someone with his left leg.
Answer: His left leg extends out with force as he kicks something or someone. His arms are held steady at his sides. His torso is slightly twisted. His buttocks and waist muscles are contracted. His neck remains neutral.
These examples are essential as they are critical in aiding LLMs in maintaining the structural stability of their output results. This design guarantees the stable quality of the fine-grained descriptions while re-annotating thousands of textual descriptions from HumanML3D and KIT.
@@ 3.2 Diffusion Model for Motion Generation
The diffusion model consists of the forward process and the reverse process, both of which follow the Markov chain. The forward process involves adding noise. The input is the original motion $x_{0}^{1:n}$ from the data distribution, and the output is the motion $x_{t}^{1:n}$ with adding Gaussian noise $t$ times. When enough noise is added, the motion $x_{T}^{1:n}$ can approach the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process aims to reduce the noise in the Gaussian noise $x_{T}^{1:n}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In the denoising process, at diffusion step $t$, a portion of the noise is eliminated, resulting in a less noisy motion $x_{t-1}^{1:n}$. This step is repeated iteratively until the noise is completely removed, generating a clean motion $\hat{x}_{0}^{1:n}$.
@@ 3.2.1 Network
we adopt a simple transformer [44] encoder architecture to implement our network $G$. Unlike the conventional diffusion model mentioned above, we follow [36] and predict the clean motion $\hat{x}_{0}^{1:n}$ instead of predicting the noise added in each time-step. The input of $G$ is the noised motion $x_{t}^{1:n}$ obtained by adding noise $t$ times to the original motion $x_{0}^{1:n}$. The noised motion $x_{t}^{1:n}$, together with the text condition tokens $GL$, $PT_{1:k}$ and the time-step $t$, are inputted to the transformer encoder, resulting in the clean motion $\hat{x}_{0}^{1:n}$. One of the reasons for directly predicting the clean motion in each time-step of the diffusion model is to incorporate human geometric losses during the training of the network, making the generated human motions more natural. For each sampling step $t$, from $T$ to 1, our model predicts the clean motion $\hat{x}_{0}^{1:n}$, and then adds noise back to $x_{t-1}^{1:n}$. After $T$ iterations, the final clean motion $\hat{x}_{0}^{1:n}$ is obtained. This form of diffusion model has become commonly adopted, as do we.